OpenStudy (anonymous):

please help me, i got final exam tomorrow integrate : [(2+arctan 3x)^(1/2) / 1+9x^2

OpenStudy (anonymous):

Let\[3x =\theta\]

OpenStudy (amistre64):

let 3x=theta?

OpenStudy (amistre64):

the bottom is the derivative of the top :)

OpenStudy (amistre64):

well, the top innnards

OpenStudy (anonymous):

May be you have an easier way.

OpenStudy (amistre64):

trig sub it lol.. no really u = 2+tan^-1(3x) du = 3/(1+9x^2) dx dx = (1+9x^2)/3

OpenStudy (amistre64):

(1+9x^2) u^(1/2) --------------- (1+9x^2) 3 (1/3) [S] u^(1/2) du

OpenStudy (anonymous):

Good, my way was much more labor.

OpenStudy (amistre64):

lol.... i like laborious :)